OpenStudy (anonymous):

How do you graph x-3=y?

OpenStudy (anonymous):

Help Please?

OpenStudy (anonymous):

Draw a Cartesian Plane (x and y axis) with an appropriate scale. Then plots the points according to the equation. e.g. x-3=y means that every y-coordinate is equal to the x-coordinate take three. So, when x = 3, (3)-3=y y=0 Therefore when x = 3, y = 0 or (3, 0) Do this for a couple of points and then since the equation is linear draw a line through the points.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Plot+x-3%3Dy